What is the exact value of :
arcsin (-sqrt(3)/4)
cot(tan inverse 5)
cos inverse (sin5pi/4)
sine(sine inverse 2/3)

Answer by stanbon(75887)

arcsin (-sqrt(3)/4) = -25.66 degrees or +334.3411 degrees or 205.6589 degrees
-------------------------
cot(tan inverse 5) = 1/5
------------------------------
cos inverse (sin5pi/4)
sin(5pi/4) = -0.7071
---
cos^-1(-0.7071) = 135 degrees
-------------------------------
sine (sine inverse 2/3) = 2/3
===============================
Cheers,
Stan H.
================